# On Lepton and Baryon Numbers as Local Gauge Symmetries

### v1

@@ II Local Lepton Number
Notice that since ${\mathcal{A}}(SU(2)_{L}^{2}U(1)_{\ell})\neq 0$ one needs at least one fermionic representation transforming non-trivially under $SU(2)_{L}$. Different solutions have been proposed for anomaly cancellation in this context: a) Vector-like leptons Fileviez Perez and Wise (2011); Duerr et al. (2013), b) Four fermionic representations Fileviez Perez et al. (2014). If one stick to the minimal number of fields needed to understand the spontaneous breaking of total lepton number the neutrinos are Dirac fermions in these theories because the total lepton number is broken by three units. The case of Majorana neutrinos has been also discussed in Refs. Debnath and Fileviez Perez (2023); Debnath et al. (2024).
@@
    with $\tilde{M}_{D}^{\nu}=y_{\nu}v_{0}N_{R}/\sqrt{2}$ and $M_{\rho}=\lambda_{\rho}v_{S}/\sqrt{2}$. Here $v_{0}=246$ GeV is the vacuum expectation value of the SM Higgs field and $M_{R}$ is the mass of the massive right-handed neutrino. Here we use the convention where only the third right-handed field is massive. In the above equation we included the mass for $\rho_{L}^{0}$ and the mixing term between $\nu_{L}$ and $\rho_{L}^{0}$.
@@
    Here $m_{D}^{i}=y^{i}v_{0}v_{\phi}/4\Lambda$. Notice that $m_{\nu}^{ij}$ has rank two. Therefore, this theory predicts a 3+2 light neutrino sector. This theory does not predict the numerical values of the coefficients in the above mass matrix, but one expects that the two extra sterile neutrinos can have mass below or at the eV scale.
@@
    The mixing between the SM charged leptons and the new charged lepton is very small since $v_{\phi},v_{S}\gg v_{0}$. This simple theory predicts lepton flavour violation mediated by the the interactions in Eq.(3) that can give rise to lepton flavour violating processes such as $\mu\to e\gamma$. We will discuss the predictions for these processes in a future publication. The charged components in $\rho_{L}$ mix with the SM charged leptons via the dimension five operator in Eq.(3).
  • Proton Decay: In this theory proton decay can be mediated by the dimension ten operators [^1] Thanks to Mark B. Wise for this comment.:
    $\frac{c_{\ell}}{\Lambda^{6}}q_{L}q_{L}q_{L}\ell_{L}\phi^{4}+\frac{c_{R}}{\Lambda^{6}}u_{R}u_{R}d_{R}e_{R}\phi^{4}.$ (15)
    Using $v_{\phi}\sim 1$ TeV one finds that $\Lambda>10^{4}$ TeV in order to satisfy the experimental bounds on the proton decay lifetime. For a review on proton decay see Ref. Nath and Fileviez Perez (2007).
@@ III Local Baryon Number
$SU(3)_{C}\otimes SU(2_{L}\otimes U(1)_{Y}\otimes U(1)_{\ell}\otimes U(1)_{B}.$
Using the SM quark fields: $q_{L}\sim({\bf{3}},{\bf{2}},1/6,1/3)$, $u_{R}\sim({\bf{3}},{\bf{1}},2/3,1/3)$ and $d_{R}\sim({\bf{3}},{\bf{1}},-1/3,1/3)$, one can estimate the baryonic gauge anomalies:
@@ IV Summary
In this article we have discussed the simplest theory based on local total lepton number where the gauge anomalies are cancelled with only four fermions. The neutrinos are predicted to be Majorana fermions. The neutrino masses are generated through a double seesaw mechanism predicting a light $3+2$ neutrino sector, where one expects that the two extra sterile neutrinos can have mass below or at the eV scale. We have shown that using the same extra fermionic fields one can cancel all baryonic anomalies and define the simplest theory where the total lepton and baryon numbers can be spontaneously broken at the low scale because the proton is stable and other baryon number violating processes are highly suppressed.
Acknowledgments: I would like to thank M. B. Wise for discussions. This work has been supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0024160.

### v2

@@ II Local Lepton Number
Notice that since ${\mathcal{A}}(SU(2)_{L}^{2}U(1)_{\ell})\neq 0$ one needs at least one fermionic representation transforming non-trivially under $SU(2)_{L}$. Different solutions have been proposed for anomaly cancellation in this context: a) Vector-like leptons Fileviez Perez and Wise (2011); Duerr et al. (2013), b) Four fermionic representations Fileviez Perez et al. (2014). If one stick to the minimal number of fields needed to understand the spontaneous breaking of total lepton number, the neutrinos are Dirac fermions in these theories because the total lepton number is broken by three units. The case of Majorana neutrinos has been also discussed in Refs. Debnath and Fileviez Perez (2023); Debnath et al. (2024).
@@
    with $\tilde{M}_{D}^{\nu}=y_{\nu}v_{0}N_{R}/\sqrt{2}$ and $M_{\rho}=\lambda_{\rho}v_{S}/\sqrt{2}$. Here $v_{0}=246$ GeV is the vacuum expectation value of the SM Higgs field and $M_{R}$ is the mass of the massive right-handed neutrino. Here we use the convention where only the third right-handed neutrino field is massive. In the above equation we included the mass for $\rho_{L}^{0}$ and the mixing term between $\nu_{L}$ and $\rho_{L}^{0}$.
@@
    Here $m_{D}^{i}=y^{i}v_{0}v_{\phi}/4\Lambda$. Notice that $m_{\nu}^{ij}$ has rank two. Therefore, this theory predicts a 3+2 light neutrino sector. This theory does not predict the numerical values of the coefficients in the above mass matrix, but one expects that the two extra sterile neutrinos can have masses smaller than the SM neutrino masses with small mixing angles. This is possible when $(\tilde{M}_{D}^{\nu})^{i\alpha}$ is very small. Notice that since $m_{\nu}^{ij}$ has rank two, one of the SM neutrinos must be massless. We will study in detail the cosmological bounds in this model in a future publication.
    In this discussion, we did not include the dimension five operators such as $H^{\dagger}\rho_{L}\chi_{L}H/\Lambda$ because generically it provides a small corrections to the $\rho_{L}$ and $\chi_{L}$ masses. One could have the dimension six operator: $c_{R}\nu_{R}\nu_{R}\phi^{2}S^{*}/\Lambda^{2}$ that can change the predictions for the right-handed neutrino masses. Assuming $c_{R}\sim 1$, $v_{\phi}\sim v_{S}\sim 1$ TeV and $\Lambda>10^{5}$ TeV (see discussion about proton decay), one finds that $\delta M_{R}<50$ eV. Notice that if this operator is present one changes the tree-level predictions for the right-handed neutrino masses above, but still two right-handed neutrinos are very light.
@@
    The mixing between the SM charged leptons and the new charged lepton is very small since $v_{\phi},v_{S}\gg v_{0}$. This simple theory predicts lepton flavour violation mediated by the interactions in Eq.(3) that can give rise to lepton flavour violating processes such as $\mu\to e\gamma$. We will discuss the predictions for these processes in a future publication. The charged components in $\rho_{L}$ mix with the SM charged leptons via the dimension five operator in Eq.(3).
  • Proton Decay: In this theory proton decay can be mediated by the dimension nine operators [^1] Thanks to C. Murgui and the referee for this comment.. For example, one can have these operators:
    $\frac{c_{\ell}}{\Lambda^{6}}q_{L}q_{L}q_{L}\ell_{L}S^{*}(\phi^{*})^{2}+\frac{c_{e}}{\Lambda^{6}}u_{R}u_{R}d_{R}e_{R}S^{*}(\phi^{*})^{2}.$ (15)
    Using $c_{\ell}\sim 1$, $c_{e}\sim 1$, $v_{\phi}\sim v_{S}\sim 1$ TeV, one finds that $\Lambda>10^{5}$ TeV in order to satisfy the experimental bounds on the proton decay lifetime. For a review on proton decay see Ref. Nath and Fileviez Perez (2007). Notice that these operators are not allowed in the full theory that will be discuss in the next section, where the baryon number is also gauged.
In this section, we have assumed the scalar field $\phi$ acquires a vacuum expectation value generating the mixings between the new fields and the SM leptons. This allows us to generate the neutrino masses and as we will discuss in the next section, we need only two scalar fields, $S$ and $\phi$, to define a realistic theory where the total lepton and baryon numbers are spontaneously broken. If $\phi$ does not have a vacuum expectation value, the model can be realistic because the new fields can decay into the SM fields and one of the new neutral fields. In this case, the lightest field between $\phi$, $\chi_{L}$ and $\rho_{L}^{0}$ can be a dark matter candidate.
@@ III Local Baryon Number
$SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y}\otimes U(1)_{\ell}\otimes U(1)_{B}.$
Using the SM quark fields: $q_{L}\sim({\bf{3}},{\bf{2}},1/6,0,1/3)$, $u_{R}\sim({\bf{3}},{\bf{1}},2/3,0,1/3)$ and $d_{R}\sim({\bf{3}},{\bf{1}},-1/3,0,1/3)$, one can estimate the baryonic gauge anomalies:
@@
In this section, we have discussed the full theory where the baryon and total lepton numbers are local gauge symmetries. One can consider the case where only the baryon number is a local gauge symmetry and cancel all the baryonic anomalies with the four extra fermions, $\Psi_{L},\Psi_{R},\rho_{L}$ and $\chi_{L}$, listed above. In this case, the $\phi$ scalar field is still needed to allow the new $\Psi$-fields to decay. Now, in the scenario where $\phi$ does not acquire a vacuum expectation value, the lightest field between $\phi$, $\chi_{L}$ and $\rho_{L}^{0}$ is stable and it can be a cold dark matter candidate. In this case, one has an accidental discrete symmetry protecting the stability of the lightest field, i.e. $\mathcal{Z}_{2}$: $\phi\to-\phi$, $\chi_{L}\to-\chi_{L}$, $\rho_{L}\to-\rho_{L}$, $\Psi_{L}\to-\Psi_{L}$, and $\Psi_{R}\to-\Psi_{R}$. This theory will be investigated in a future publication.
@@ IV Summary
In this article, we have discussed the simplest theory based on local total lepton number where the gauge anomalies are cancelled with only four fermions plus the right-handed neutrinos needed for the seesaw mechanism. The neutrinos are predicted to be Majorana fermions. The neutrino masses are generated through a double seesaw mechanism predicting a light $3+2$ neutrino sector, where one expects that the two extra sterile neutrinos can have mass below or at the eV scale. We have shown that using the same extra fermionic fields one can cancel all baryonic anomalies and define the simplest theory where the total lepton and baryon numbers can be spontaneously broken at the low scale. In this theory, the proton is stable and other baryon number violating processes are highly suppressed. We also briefly discussed the scenarios where only one symmetry, baryon or total lepton number, is gauged. We discussed the possibility to have a dark matter candidate in these scenarios.
Acknowledgments: I would like to thank C. Murgui and M. B. Wise for discussions. This work has been supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0024160. I would like to thank the referees for very useful comments.